1. Calculate the molar mass of Calcium Acetate, Ca(C2H302)2 Show your work, include units
Answer #1

Molar mass of Ca = 40.07g/mol

Molar Mass of C = 12.01 g/mol

Molar Mass of H = 1.00 g/mol

Molar mass of O = 16g/mol

In Calcium Acetate, Number of

a) Ca = 1

b) C= 2*2 =4

c) H = 3*2=6

d)O = 2*2 = 4

So Molar mass of Calcium Acetate = (1*Molar Mass of Ca) + (4*Molar Mass of C)+(6*Molar Mass of H)+(4*Molar Mass of O)

= (1*40g/mol)+(4*12.01g/mol)+(6*1.00g/mol)+(4*16g/mol)

=158.04g/mol (Answer)